Tags:
create new tag
, view all tags

Calorimeter Electronics Status


Below are compiled useful information addressing the current status of the calorimeter Electronics during the installation and the commissioning phase toward the final setup. The expected final readout setup is described here External link mark
and is summarized in the dedicated LHCbCaloReadout wiki page.



ECAL/HCAL

Masked Channels

Crate Slot/Feb Channel Input Observation Mask date Unmask date
22 (HCAL) 3/3 31 Noisy channel with sometimes high signals March 17th, 2010 April 8th, 2010
11 (ECAL) 8/6 5 Noisy channel April 8th, 2010 May 20th, 2010
24 (HCAL) 11/8 7 Noisy channel July 24th, 2010 January 19th, 2011
22 (HCAL) 8/6 11 Noisy channel July 24th, 2010 January 19th, 2011

FEB

Spares status

FEB Status Poignées Observations
006 Good Noires Spare
007 Good Noires Spare
018 Good   Spare - TVB specific
049 Good   Spare
062 Good   Spare
110 Good   Spare - Screew broken in front panel connector - Used for crate clock phase monitoring
139 Good   Spare
158 Good   Spare - TVB specific
173 Good  

Missing bit -> BGA replacement. Afterwards, ECS problems appeared (March 29th, 2010)

BGA has to be replaced again (could not read the checksum with the Silicon explorer). During the operation some contacts have been pulled out and the board had to be fixed before a new fpga could be soldered. Finally the board was repaired and works.

191 Good   Spare
216 Faulty  

Seq PGA problem - BGA warms up - jtag scan chain failure - BGA replacement

Seq PGA replaced on January 19, 2011. Problem remains.

223 Good   Spare
236 Faulty   Trigger PGA problem -> BGA to replace ? can be used as a pin-diode - Sent to LAL from crate timing studies.
253 Faulty   A channel is noisy. Board removed from the calorimeter.

Spare installed back on the detector

FEB Status Date Observations
060 Good    
253 Good January 18th, 2011 A channel was noisy at the pit -> Could not observe it in test bench
191 Good February 3rd, 2011  

"Burnt" FEB

86 93 81 170 38 95 85 39

Doubtful connections

bad_connections080410.xls

Specific FEB

FEB with fugitive problem

FEB Status Observation
006 Spare Channel input 13 does not work properly - Remove from the calorimeter on March 15th, 2010 -> Could not observe it in test bench
007 Spare SPECS problem have been observed at the pit and in test area. But problem diasappeared during the latest debugging (March 09, 2010).
139 Spare A channel was noisy at the pit (channel 31) -> Could not observe it in test bench

Weak FEB HCAL ->TVB trigger signal

FEB 018 158

ECAL only

FEB Status Observation
272 Installed in ECAL Problem with a trigger (HCAL connection) serialiser

Not functional FEBs

FEB Observation
035 "Burnt" board used to define temperature profil for BGA replacement -> destroyed

History of debugging/reparations

FEB

Pending

Issue

Fixed on Intervention
060   March 23rd, 2010 Missing bit (value 8) on third channel BGA 2 - BGA replaced
062   March 29th, 2010 Missing bit on pga 3/adc 0. BGA replaced. No improvement. Board to be debugged again.
    January 19th, 2011 Missing bit -> BGA FE3 replacement after soldering a wire through the PCB. Problem solved.
110   January 19th, 2011  
140   March 23rd, 2010 Delay Chip replacement
173 x March 29th, 2010 Missing bit on pga 1. NGA replaced. PGA no functional anymore (no ECS).
    January 19th, 2011 BGA has been removed but is still at the electronics service.
191   January 19th, 2011 Delay chip seqtrig was dead. Replaced.
216 x January 19th, 2011 Seq was changed (not functioning and very hot). After replacement, problem remains.
223   March 23rd, 2010 ECS problem on FEPGA 6 - BGA replaced

CROC

Spares

CROC Status Observation
  Good Spare but located at LAL - No shielding (removed for activation studies)
  Good Spare
  Good Spare
  Good Spare
  Good Spare
  Good Spare

Cabling

Due to backplane problem the following FEB are currently moved with respect to their expected location.
Ecal crate #16 : FEB#12 and FEB#13 moved FROM slot 17 and 18 TO slot 18 and 19
Ecal crate #15 : FEB#13 moved FROM slot 18 TO slot 19 WARNING this slot is to be used by one LEDTSB
Hcal crate #22 : FEB#14 moved FROM slot 19 TO slot 20 NO problem as it's a PIN readout FEB

Comment : backplane from Ecal crate#15 and #16 need to be repared. Moving FEB affects L0Calo.

PRS/SPD

Cabling

TELL1

Tell1 boards and crates

Trigger

Trigger Validation Boards

* Scan of TVB inputs:

Board Scan of HCAL Inputs Scan of ECAL Inputs
0 HCAL0 PatternA, HCAL0 PatternB Crate 8 slot 0-7
HCAL1 PatternA, HCAL1 PatternB
HCAL2 PatternA, HCAL2 PatternB
HCAL3 PatternA, HCAL3 PatternB
1 HCAL0 PatternA, HCAL0 PatternB Crate 8 slot 11-18
HCAL1 PatternA, HCAL1 PatternB
HCAL2 PatternA, HCAL2 PatternB
2 HCAL0 PatternA, HCAL0 PatternB Crate 9 slot 0-7
HCAL1 PatternA, HCAL1 PatternB
HCAL2 PatternA, HCAL2 PatternB
HCAL3 PatternA, HCAL3 PatternB
3 HCAL0 PatternA, HCAL0 PatternB Crate 9 slot 11-18
HCAL1 PatternA, HCAL1 PatternB
HCAL2 PatternA, HCAL2 PatternB
4 HCAL0 PatternA, HCAL0 PatternB Crate 10 slot 0-7
HCAL1 PatternA, HCAL1 PatternB
HCAL2 PatternA, HCAL2 PatternB
HCAL3 PatternA, HCAL3 PatternB
5 HCAL0 PatternA, HCAL0 PatternB Crate 10 slot 11-18
HCAL1 PatternA, HCAL1 PatternB
HCAL2 PatternA, HCAL2 PatternB
HCAL3 PatternA, HCAL3 PatternB
6 HCAL0 PatternA, HCAL0 PatternB Crate 11 slot 0-7
HCAL1 PatternA, HCAL1 PatternB
HCAL2 PatternA, HCAL2 PatternB
7 HCAL0 PatternA, HCAL0 PatternB Crate 11 slot 11-18
HCAL1 PatternA, HCAL1 PatternB
HCAL2 PatternA, HCAL2 PatternB
8 HCAL0 PatternA, HCAL0 PatternB Crate 12 slot 0-7
HCAL1 PatternA, HCAL1 PatternB
HCAL2 PatternA, HCAL2 PatternB
9 HCAL0 PatternA, HCAL0 PatternB Crate 12 slot 11-18
HCAL1 PatternA, HCAL1 PatternB
HCAL3 PatternA, HCAL3 PatternB
10 HCAL0 PatternA, HCAL0 PatternB Crate 13 slot 0-7
11 HCAL0 PatternA, HCAL0 PatternB Crate 13 slot 11-18
HCAL1 PatternA, HCAL1 PatternB
12 HCAL0 PatternA, HCAL0 PatternB Crate 14 slot 0-7
HCAL1 PatternA, HCAL1 PatternB
13 HCAL0 PatternA, HCAL0 PatternB Crate 14 slot 11-18

-- FredericMachefert - 15 March 2010

Topic attachments
I Attachment Action Size Date Who Comment
Microsoft Excel Spreadsheetxls bad_connections080410.xls manage 9.0 K 2010-04-08 - 10:05 FredericMachefert  
Microsoft Excel Spreadsheetxls bad_connections150310.xls manage 9.0 K 2010-03-15 - 23:55 FredericMachefert Bad ECAL/HCAL connections (fixed by pushing the cable)
Topic revision: r25 - 2011-09-02 - FredericMachefert
 

TWIKI.NET
This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2022 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback